OpenStudy (anonymous):

Can someone hlep me with this math problem? serie : an=(alpha*n)/(n+1), n>=1 (alpha is from R-real number). I must determine if the series is monotone and bouded thank you very much

OpenStudy (anonymous):

i would try with \(\alpha =1\) first

OpenStudy (anonymous):

in this case you would get \(a_n=\frac{n}{n+1}\) which is bounded below by \(1\) since \(\frac{n}{n+1}<1\) for all \(n\) and \(\lim_{n\to \infty}\frac{n}{n+1}=1\)

OpenStudy (anonymous):

i meant "bounded above by one" not below!

OpenStudy (anonymous):

it is monotone increasing as you can check by taking the derivative and seeing that is always positive

OpenStudy (anonymous):

then try the general case with \(\alpha\)

OpenStudy (anonymous):

yes alpha* n/(n+1) limi\[\lim_{x \rightarrow \infty}{n/n+1}\] \[n/n(1+1/n)=1/(1+1/n)\] take limit to infinity \[1/(1+0)=1\] \[\alpha*1=alpha\]

OpenStudy (anonymous):

the last term is alpha

OpenStudy (anonymous):

the serries os monotone

OpenStudy (anonymous):

the serries is also bounded below becaise it says n>1 soo first term is 1

OpenStudy (anonymous):

plz give me medal if i helped